OpenStudy (anonymous):

please help, find the slope (m) and the y intercept (b) of each line number 1: y=x+1 number 2: y=1/2x-2 number 3: 2x+y=3 PLEASE help and show work? i have no clue

OpenStudy (amistre64):

what is the slope intercept form of a line in general?

OpenStudy (anonymous):

i dont know.. this is all it says on my summer math packet

OpenStudy (immanuelv):

slope for #1 is 1

OpenStudy (anonymous):

the slope is whatever is in front of the x so the first is 1 the second is 1/2 and the third u need to make y=mx+b so y=-2x+3 so the last is -3

OpenStudy (amistre64):

im pretty sure they give you more than that in the material ....

OpenStudy (anonymous):

the y intercept is the last number so 1 -2 3

OpenStudy (amistre64):

how would they expect you to answer the questions if they dont provide the basis for it?

OpenStudy (anonymous):

they didnt. its my high school summer math packet and thats all it says on the directions :(

OpenStudy (anonymous):

in y=mx+b the m is slope and b is y intercept

OpenStudy (amistre64):

well, if you want, we could cycle thru the many "forms" that an equation of a line makes ...

OpenStudy (anonymous):

so its just that easy? i just need to look at the equation not even try to do any work?

OpenStudy (immanuelv):

anything infront of "X" is the slope...

OpenStudy (amistre64):

the work is to recognize the general form of the equation and pick out the important parts :)

OpenStudy (amistre64):

"anything in front of x" is not quite right; you have to have the proper form first

OpenStudy (amistre64):

3x =4y -7 ; the slope is not 3 ....

OpenStudy (anonymous):

no because its not y=mx+b that's standard form

OpenStudy (amistre64):

exactly :)

OpenStudy (anonymous):

so whats the right answer? is superjew right?

OpenStudy (anonymous):

yes im right

OpenStudy (anonymous):

ok thank you so much! :)